# Emulating 2D Materials with Magnons

### v1

@@ V.1 Inversion-broken crystals
In graphene, the masslessness of Dirac electrons is protected in part by inversion symmetry. Hexagonal boron nitride (h-BN), another 2D compound, is isostructural but with different atoms on each honeycomb sublattice. This breaks the inversion symmetry of the unit cell, and the Dirac points gap (for this reason, these symmetry-breaking terms are sometimes called mass terms). As a result, h-BN is an insulator. This is shown schematically in Fig. 6(b). Even stacking graphene on h-BN has this result because of the small amount of symmetry breaking [giovannetti_substrate-induced_2007, xue_scanning_2011]. This principle is powerful because, unless the system is fine-tuned, any method of breaking inversion symmetry should result in a gap.
@@
In graphene-like systems, these symmetry-breaking mass terms do more than simply open a gap. For small mass terms, opening such a gap generates sharply peaked Berry curvature near the valleys $\mathbf{K}$ and $\mathbf{K^{\prime}}$. Integrating the Berry curvature over the vicinity of a single valley yields an approximately quantized value, often referred to as a valley Chern number (VCN) [zhang_valley_2013]. The two valleys carry opposite VCNs, so the total Chern number over the full Brillouin zone is zero. Nevertheless, a change in VCN across a domain wall implies valley-polarized boundary modes [semenoff_domain_2008, jung_valley-hall_2011], which remain robust as long as intervalley scattering remains weak [zhang_valley_2013]. TB calculations in this $s,p_{x,y}$ system show that symmetry breaking has this effect for both the $s$-band gaps and the $p$-band gaps. At a single valley, the sign of the Berry curvature is opposite above and below the gap for both $s$ and $p$ bands, but notably has a different sign in each ($s,p$) case. Upon changing the sign of the symmetry-breaking terms $\Delta\varepsilon_{s}$ and $\Delta\varepsilon_{p}$, the sign of curvature flips (see supplemental Sec. LABEL:Supp-sec:S_BerryCurvature), schematically illustrated in Fig. 7(a).
This magnetic thin film geometry, unlike vdW systems, is continuously tunable between distinct gapped phases. Partially inspired by the Jackiw-Rebbi model [jackiw_solitons_1976], we investigate a boundary between these gapped phases. Figure 7(b) shows a gradual phase boundary in which the unit cell changes shape over a few lattice spacings, making a smooth transition between the two phases.
@@
In our TB model, as we increase mass terms, the Berry curvature is no longer sharply localized near the valleys and the VCN loses its quantization (see supplemental Sec. LABEL:Supp-sec:S_BerryCurvature). This behavior is common in realistic systems [lee_gapped_2022, zhu_design_2018, lee_quantum_2020], where the absolute value of the VCN is smaller than its quantized value in the small-gap limit and varies smoothly with system parameters. However, since the bulk gap does not close as we vary mass parameters, the boundary modes we observe are adiabatically connected to the boundary modes of the small-gap limit. They remain (for example, in our simulations) visible well beyond the regime where the VCN is sharply quantized. This phenomenon of boundary-localized states has been seen before in many other graphene-like systems [lu_observation_2017, yang_acoustic_2018, wang_extended_2022, noh_observation_2018, zhang_achieving_2018, wang_ultracompact_2022, funayama_quantum_2024], but occurs here for both the $s$ and $p$ bands. Because the $\Delta\varepsilon$ terms affect the $s$ and $p$ gaps in different ways, the corresponding magnon boundary modes bridge the gaps with opposite chiralities: in the lower gap, $\mathbf{K}$($\mathbf{K^{\prime}}$)-valley magnons propagate to the left (right), while in the upper gap, $\mathbf{K}$($\mathbf{K^{\prime}}$)-valley magnons propagate to the left (right). The direction of this valley-polarized magnon current can be reversed by changing the sign of the symmetry-breaking terms (see supplemental Sec. LABEL:Supp-sec:S_InvertedBoundary). By calculating the VCNs of the bulk bands in the small-gap limit, we thus identify these boundary states as the magnonic analog of quantum valley Hall (QVH) edge states[yao_edge_2009, zhang_valley_2013], as illustrated in Fig. 7(e). This is a more accessible realization of the type of state proposed for 2D honeycomb magnets with controllable staggered anisotropies [hidalgo-sacoto_magnon_2020]. For a more complete discussion including Berry curvature maps, see the supplemental Sec. LABEL:Supp-sec:S_BerryCurvature.
@@ V.2 Topological Magnon Waveguiding
In the same geometry as before, shown again in Fig. 8(a), an excitation is applied inside a localized region. At the in-gap frequency of $2.7\text{ GHz}$, spin waves are launched only along the phase boundary, as shown in Fig. 8(b). Because of the valley-specific topological origin of these modes, left- and right- moving modes are oppositely valley polarized – an interesting feature because valley polarization is occasionally thought of as an information-carrying degree of freedom in discussions of next-generation information technologies [schaibley_valleytronics_2016]. The transition between the insulating phases can be reduced in width, and the states persist, even following sharp turns made by the phase boundary. This phenomenon occurs for both $s$ and $p$ boundary modes [see Figs. 8(c) and (d), respectively]. In both cases, magnon transport remains completely localized to the boundary after $500\text{ ns}$ of continuous excitation. It is worth noting a few limitations here. Geometry does not only affect the shape of the band structure; there can be a constant frequency offset as a result of the demagnetizing field. This is the same reason $B_{\text{ext}}$ is varied for the different geometries in Fig. 2. When geometry is varied over space, this varying frequency offset can result in a spectral overlap of the localized mode with other bulk modes. In this case, a large-area excitation like that in Fig. 8(a) would excite both boundary and bulk modes. In other words, the appealing boundary-specific transport demonstrated by Fig. 8 may not exist in all geometries with phase boundaries. Also note that these modes are not protected against backscattering, making them fundamentally different from more famous examples of topological edge states in metamaterials, in [shindou_topological_2013, hafezi_imaging_2013] for example. Scattering between valley modes caused by a point defect is demonstrated in Fig. 8(e-g) for both $s$-gap and $p$-gap modes, returning differing reflection amplitudes for each case. This is discussed in more detail in the supplemental Sec. LABEL:Supp-sec:S_Backscattering.
@@ V.3 Isolated defect modes
In a phenomenon similar to spin wave ”edge modes,” [maranville_characterization_2006, mcmichael_edge_2006] A single defect has a localized mode which lies below the bulk uniform precession mode. Two such defect modes couple when placed in adjacent sites. Using the TB interpretation, the level should split into symmetric and antisymmetric superpositions of the original modes, separated in frequency by twice the coupling parameter $\Delta$. This is apparently a good interpretation, as evidenced by the mode profiles in Fig. 9(b). A single-defect mode at $1.47\text{ GHz}$ splits into a symmetric mode at $1.38\text{ GHz}$ and an antisymmetric mode at $1.56\text{ GHz}$ (for details, see supplemental Sec. LABEL:Supp-sec:S_Defects). It happens that this whole level structure appears below the continuum, which implies that these defect-localized magnons may be interacted with separately from propagating waves, suggesting further application in magnonics. For the chosen Gilbert damping parameter of $\alpha=10^{-4}$, these modes have quality factors $\approx 4100$.
@@ VI Other TB-like systems
Figure 10: Other geometries which can be thought of as TB-like, using 15-nm YIG films with a lattice spacing a = 300 nm and $B_{\text{ext}}||\hat{z}=185\text{ mT}$. (a) The square anti-dot lattice, and (b) its TB model, using (c) basis orbitals similar to those used in the hexagonal anti-dot model. Some inconsistencies are due to hybridization with higher-order modes like $d$-orbitals (see supplemental Sec. LABEL:Supp-sec:S_OtherTB), which are not included for simplicity. (d) Shows the artificial kagome lattice and (e) its TB model using only $s$ orbitals on kagome sites.
The magnonic structure of the perpendicularly magnetized square anti-dot lattice [see Fig. 10(a)] happens to be analyzable using the same approach. A TB model with $s$ and $p$ orbitals in the middle of the unit cell and $s$ orbitals between these sites can approximate the band structure with the right fitting parameters. Its spectrum is plotted in Fig. 10(b), using the basis orbitals in Fig. 10(c). This square lattice is apparently another case of emergent ”atomic orbitals” (see supplemental Sec. LABEL:Supp-sec:S_OtherTB for details). There is a simpler (albeit less experimentally feasible) approach to engineer magnonic band structures. Figure 10 also shows an artificial kagome lattice, where it is assumed that disks will host $s$-like modes, so the disks are arranged to mimic the kagome lattice [see Fig. 10(d)]. In this case, a simple model of $s$-orbitals on the kagome lattice can approximate the band structure [see Fig. 10(e)]. This is similar to results previously reported for magnetic rods embedded in a magnetic matrix [yang_flatbands_2025, liang_dirac_2024]. In this case, the relative decrease in momentum space resolution is due to the fact that the unit cell must be simulated with greater resolution, so the whole simulation field has fewer unit cells for computationally practical reasons.
@@ VII Experimental feasibility
The YIG anti-dot lattices studied in this work may be fabricated by standard electron beam lithography procedures including a hard mask and ion etch step, or by focused ion beam (FIB) milling for the fabrication of single devices. There is already a significant amount of work on these structures [neusser_anisotropic_2010, ulrichs_magnonic_2010, bali_high-symmetry_2012, gros_phase_2021, wang_observation_2023], albeit with somewhat different dimensions. Preliminary work indicates that FIB can yield devices with sufficient scale and tolerances. In the supplemental material (Sec. LABEL:Supp-sec:S_Deadlayer), we verify that lattices with a thin magnetic dead layer caused by ion damage are expected to yield similar physics. We also verify that in the presence of edge roughness, flat modes, band gaps, defect-localized modes, and QVH-like modes persist with only moderate modification (Sec. LABEL:Supp-sec:S_Disorder). All simulations are conducted for a Gilbert damping parameter $\alpha=10^{-4}$ and $M_{S}=$140\text{\,}\mathrm{k}\mathrm{A}\mathrm{/}\mathrm{m}$\$, which is realistic for YIG [dubs_low_2020, li_epitaxial_2015]. Overall, the anti-dot lattice is a more realistic experimental route to these types of magnonic flat bands compared to existing proposals, which mainly focus on ferromagnetic rods embedded in a matrix[centala_compact_2023, liang_dirac_2024, yang_flatbands_2025].
@@ VIII Discussion and Conclusion
First, the existence of isotropic flat bands of Fig. 3 is notable. In the TB picture, it is trivially easy to achieve flat bands in one sense: an array of decoupled resonators has no dispersion. However, here, the modes are strongly coupled and flatness emerges from the symmetry of the basis orbitals and the geometry of the lattice. Bloch functions in this band contain singularities at band touching points, a property which sometimes awards it the title of ”topological flat band” in electronics [liu_orbital_2022, rhim_singular_2021]. Furthermore, these states may be of interest for exploring magnon-magnon interactions, particularly as analogs of interacting electronic phases on the kagome lattice. The localization property of these magnons indicates that they could easily be excited to large magnon numbers, potentially resulting in emergent modes which already have some theoretical description in the hot research area of interacting kagome flat bands. The excitation enhancement due to flat band magnons has been studied before, for example in [wang_broad-wave-vector_2024].
@@
Third, additional control over the valley degree of freedom could reasonably be exercised here through an inhomogeneous pattern that mimics a strain field. Inhomogeneous strain fields can result in an effective gauge field for Dirac point excitations, commonly called a pseudo-magnetic field [pereira_strain_2009, kang_pseudo-magnetic_2021]. The freedom associated with patterning in this system might be used to realize magnon valley filters, lenses, collimators, and (pseudo-)Landau levels for magnons [sun_magnon_2021, wei_strain-engineered_2024, jamadi_direct_2020].
Fourth, point defects such as those in Fig. 9 may be useful in magnonics because of their spatial location and spectral isolation. This implies that a carefully placed waveguide could selectively excite these quasi-uniform localized modes – making them potentially useful as a magnonic memory because spin waves will not radiate away from the defect. Additionally, the fact that a coupled two-mode structure of such defects still lies below the continuum is interesting. We envision classical spin wave analogs of effects like induced transparency and Rabi oscillation. If sufficiently long magnon lifetimes can be achieved, possibly aided by the localization of the modes[serha_ultra-long-living_2025], then this may also be an interesting perspective for designing quantum magnonic systems.
@@
Finally, because there exists similar physics in other analog systems [milicevic_type-iii_2019, gao_visualization_2024, wu_p_xy-orbital_2008, polini_artificial_2013, mangussi_multi-orbital_2020, ding_experimental_2019], it is worth noting a few novelties and benefits associated with this magnonic system. The literature tends to focus on $s$-like modes, but we show that it’s also possible to accomplish valley-specific physics with $p$-like modes, and we show that it can be understood through the same TB lens. This amplifies the potential for valley-tronics because there are two channels through which valley information may propagate. There are also benefits that are unique to magnonic systems. Most importantly, the whole band structure is tunable not just by geometry, but also by externally applied field, as was demonstrated in Fig. 6(c). This suggests that a constant microwave signal may be tuned on the fly to be blocked by band gaps, waveguided, valley- or orbitally polarized, strongly localized in flat bands, interacting with defect modes, etc. simply by varying the external field. Spin waves also are associated with a unique combination of small wavelengths and low frequencies; it is a significant result that a microwave signal can be localized to these phase boundaries by some band topology manipulation.

### v2

@@ V.1 Inversion-broken crystals
In graphene, the masslessness of Dirac electrons is protected in part by inversion symmetry. Hexagonal boron nitride (h-BN), another 2D compound, is isostructural but with different atoms on each honeycomb sublattice. This breaks the inversion symmetry of the unit cell, and the Dirac points gap (for this reason, these symmetry-breaking terms are sometimes called mass terms). As a result, h-BN is an insulator. This is shown schematically in Fig. 6(b). Even stacking graphene on h-BN has this result because of the small amount of symmetry breaking [23, 100]. This principle is powerful because, unless the system is fine-tuned, any method of breaking inversion symmetry should result in a gap.
@@
In graphene-like systems, these symmetry-breaking mass terms do more than simply open a gap. For small mass terms, opening such a gap generates sharply peaked Berry curvature near the valleys $\mathbf{K}$ and $\mathbf{K^{\prime}}$. Integrating the Berry curvature over the vicinity of a single valley yields an approximately quantized value, often referred to as a valley Chern number (VCN) [106]. The two valleys carry opposite VCNs, so the total Chern number over the full Brillouin zone is zero. Nevertheless, a change in VCN across a domain wall implies valley-polarized boundary modes [77, 37], which remain robust as long as intervalley scattering remains weak [106]. TB calculations in this $s,p_{x,y}$ system show that symmetry breaking has this effect for both the $s$-band gaps and the $p$-band gaps. At a single valley, the sign of the Berry curvature is opposite above and below the gap for both $s$ and $p$ bands, but notably has a different sign in each ($s,p$) case. Upon changing the sign of the symmetry-breaking terms $\Delta\varepsilon_{s}$ and $\Delta\varepsilon_{p}$, the sign of curvature flips (see supplemental Sec. LABEL:Supp-sec:S_BerryCurvature [1]), schematically illustrated in Fig. 7(a).
This magnetic thin film geometry, unlike vdW systems, is continuously tunable between distinct gapped phases. Partially inspired by the Jackiw-Rebbi model [34], we investigate a boundary between these gapped phases. Figure 7(b) shows a gradual phase boundary in which the unit cell changes shape over a few lattice spacings, making a smooth transition between the two phases.
@@
In our TB model, as we increase mass terms, Berry curvature is no longer localized to the valleys and the VCN loses its quantization (supplemental Sec. LABEL:Supp-sec:S_BerryCurvature [1]). This behavior is common in realistic systems [48, 108, 47], where the absolute value of the VCN is smaller than its quantized value in the small-gap limit and varies smoothly with system parameters. However, since the bulk gap does not close as we vary mass parameters, the boundary modes we observe are adiabatically connected to those in the small-gap limit. They remain (for example, in our simulations) visible well beyond the regime where the VCN is sharply quantized. This phenomenon of boundary-localized states has been seen before in many other graphene-like systems [56, 102, 92, 64, 107, 91, 19], but occurs here for both the $s$ and $p$ bands. Because the $\Delta\varepsilon$ terms affect the $s$ and $p$ gaps in different ways, the corresponding magnon boundary modes bridge the gaps with opposite chiralities: in the lower gap, $\mathbf{K}$($\mathbf{K^{\prime}}$)-valley magnons propagate to the left (right), while in the upper gap, $\mathbf{K}$($\mathbf{K^{\prime}}$)-valley magnons propagate to the left (right). The direction of this valley-polarized magnon current can be reversed by changing the sign of the symmetry-breaking terms (see supplemental Sec. LABEL:Supp-sec:S_InvertedBoundary [1]). By calculating the VCNs of the bulk bands in the small-gap limit, we thus identify these boundary states as the magnonic analog of quantum valley Hall (QVH) edge states[103, 106], as illustrated in Fig. 7(e). This is a more accessible realization of the type of state proposed for 2D honeycomb magnets with controllable staggered anisotropies [30]. For a more complete discussion including Berry curvature maps, see the supplemental Sec. LABEL:Supp-sec:S_BerryCurvature [1].
@@ V.2 Topological Magnon Waveguiding
In the same geometry as before, shown again in Fig. 8(a), an excitation is applied inside a localized region. At the in-gap frequency of $2.7\text{ GHz}$, spin waves are launched only along the phase boundary, as shown in Fig. 8(b). Because of the valley-specific topological origin of these modes, left- and right- moving modes are oppositely valley polarized – an interesting feature because valley polarization is occasionally thought of as an information-carrying degree of freedom in discussions of next-generation information technologies [75]. The transition between the insulating phases can be reduced in width, and the states persist, even following sharp turns made by the phase boundary. This phenomenon occurs for both $s$ and $p$ boundary modes [see Figs. 8(c) and (d), respectively]. In both cases, magnon transport remains completely localized to the boundary after $500\text{ ns}$ of continuous excitation. It is worth noting a few limitations here. Geometry does not only affect the shape of the band structure; there can be a constant frequency offset as a result of the demagnetizing field. This is the same reason $B_{\text{ext}}$ is varied for the different geometries in Fig. 2. When geometry is varied over space, this varying frequency offset can result in a spectral overlap of the localized mode with other bulk modes. In this case, a large-area excitation like that in Fig. 8(a) would excite both boundary and bulk modes. In other words, the appealing boundary-specific transport demonstrated by Fig. 8 may not exist in all geometries with phase boundaries. Also note that these modes are not protected against backscattering, making them fundamentally different from more famous examples of topological edge states in metamaterials, in [80, 27] for example. Scattering between valley modes caused by a point defect is demonstrated in Fig. 8(e-g) for both $s$-gap and $p$-gap modes, returning differing reflection amplitudes for each case. This is discussed in more detail in the supplemental Sec. LABEL:Supp-sec:S_Backscattering [1].
@@ V.3 Isolated defect modes
In a phenomenon similar to spin wave ”edge modes,” [58, 59] A single defect has a localized mode which lies below the bulk uniform precession mode. Two such defect modes couple when placed in adjacent sites. Using the TB interpretation, the level should split into symmetric and antisymmetric superpositions of the original modes, separated in frequency by twice the coupling parameter $\Delta$. This is apparently a good interpretation, as evidenced by the mode profiles in Fig. 9(b). A single-defect mode at $1.47\text{ GHz}$ splits into a symmetric mode at $1.38\text{ GHz}$ and an antisymmetric mode at $1.56\text{ GHz}$ (for details, see supplemental Sec. LABEL:Supp-sec:S_Defects [1]). It happens that this whole level structure appears below the continuum, which implies that these defect-localized magnons may be interacted with separately from propagating waves, suggesting further application in magnonics. For the chosen Gilbert damping parameter of $\alpha=10^{-4}$, these modes have quality factors $\approx 4100$.
@@ VI Other TB-like systems
Figure 10: Other geometries which can be thought of as TB-like, using 15-nm YIG films with a lattice spacing a = 300 nm and $B_{\text{ext}}||\hat{z}=185\text{ mT}$. (a) The square anti-dot lattice, and (b) its TB model, using (c) basis orbitals similar to those used in the hexagonal anti-dot model. Some inconsistencies are due to hybridization with higher-order modes like $d$-orbitals (see supplemental Sec. LABEL:Supp-sec:S_OtherTB [1]), which are not included for simplicity. (d) Shows the artificial kagome lattice and (e) its TB model using only $s$ orbitals on kagome sites.
The magnonic structure of the perpendicularly magnetized square anti-dot lattice [see Fig. 10(a)] happens to be analyzable using the same approach. A TB model with $s$ and $p$ orbitals in the middle of the unit cell and $s$ orbitals between these sites can approximate the band structure with the right fitting parameters. Its spectrum is plotted in Fig. 10(b), using the basis orbitals in Fig. 10(c). This square lattice is apparently another case of emergent ”atomic orbitals” (see supplemental Sec. LABEL:Supp-sec:S_OtherTB [1] for details). There is a simpler (albeit less experimentally feasible) approach to engineer magnonic band structures. Figure 10 also shows an artificial kagome lattice, where it is assumed that disks will host $s$-like modes, so the disks are arranged to mimic the kagome lattice [see Fig. 10(d)]. In this case, a simple model of $s$-orbitals on the kagome lattice can approximate the band structure [see Fig. 10(e)]. This is similar to results previously reported for magnetic rods embedded in a magnetic matrix [101, 51]. In this case, the relative decrease in momentum space resolution is due to the fact that the unit cell must be simulated with greater resolution, so the whole simulation field has fewer unit cells for computationally practical reasons.
@@ VII Experimental feasibility
The YIG anti-dot lattices studied in this work may be fabricated by standard electron beam lithography procedures including a hard mask and ion etch step, or by focused ion beam (FIB) milling for the fabrication of single devices. There is already a significant amount of work on these structures [62, 88, 2, 25, 90], albeit with somewhat different dimensions. Preliminary work indicates that FIB can yield devices with sufficient scale and tolerances. In the supplemental material Sec. LABEL:Supp-sec:S_Deadlayer [1], we verify that lattices with a thin magnetic dead layer caused by ion damage are expected to yield similar physics. We also verify that in the presence of edge roughness, flat modes, band gaps, defect-localized modes, and QVH-like modes persist with only moderate modification (Sec. LABEL:Supp-sec:S_Disorder [1]). All simulations are conducted for a Gilbert damping parameter $\alpha=10^{-4}$ and $M_{S}=$140\text{\,}\mathrm{k}\mathrm{A}\mathrm{/}\mathrm{m}$\$, which is realistic for YIG [17, 50]. Overall, the anti-dot lattice is a more realistic experimental route to these types of magnonic flat bands compared to existing proposals, which mainly focus on ferromagnetic rods embedded in a matrix[7, 51, 101].
@@ VIII Discussion and Conclusion
First, the existence of isotropic flat bands of Fig. 3 is notable. In the TB picture, it is trivially easy to achieve flat bands in one sense: an array of decoupled resonators has no dispersion. However, here, the modes are strongly coupled and flatness emerges from the symmetry of the basis orbitals and the geometry of the lattice. Bloch functions in this band contain singularities at band touching points, a property which sometimes awards it the title of ”topological flat band” in electronics [54, 73]. Furthermore, these states may be of interest for exploring magnon-magnon interactions, particularly as analogs of interacting electronic phases on the kagome lattice. The localization property of these magnons indicates that they could easily be excited to large magnon numbers, potentially resulting in emergent modes which already have some theoretical description in the hot research area of interacting kagome flat bands. The excitation enhancement due to flat band magnons has been studied before, for example in [93].
@@
Third, additional control over the valley degree of freedom could reasonably be exercised here through an inhomogeneous pattern that mimics a strain field. Inhomogeneous strain fields can result in an effective gauge field for Dirac point excitations, commonly called a pseudo-magnetic field [68, 39]. The freedom associated with patterning in this system might be used to realize magnon valley filters, lenses, collimators, and (pseudo-)Landau levels for magnons [84, 97, 35].
Fourth, point defects such as those in Fig. 9 may be useful in magnonics because of their spatial localization and spectral isolation. This implies that a carefully placed waveguide could selectively excite these quasi-uniform localized modes – making them potentially useful as a magnonic memory because spin waves will not radiate away from the defect. Additionally, the fact that a coupled two-mode structure of such defects still lies below the continuum is interesting. We envision classical spin wave analogs of effects like induced transparency and Rabi oscillation. If sufficiently long magnon lifetimes can be achieved, possibly aided by the localization of the modes[79], then this may also be an interesting perspective for designing quantum magnonic systems.
@@
Finally, because there exists similar physics in other analog systems [60, 20, 99, 70, 57, 16], it is worth noting a few novelties and benefits associated with this magnonic system. The literature tends to focus on $s$-like modes, but we show that it’s also possible to accomplish valley-specific physics with $p$-like modes, and we show that it can be understood through the same TB lens. This amplifies the potential for valley-tronics because there are two channels through which valley information may propagate. There are also benefits that are unique to magnonic systems. Most importantly, the whole band structure is tunable not just by geometry, but also by externally applied field, as was demonstrated in Fig. 6(c). This suggests that a constant microwave signal may be tuned on the fly to be blocked by band gaps, waveguided, valley- or orbitally polarized, strongly localized in flat bands, interacting with defect modes, etc. simply by varying the external field. Spin waves also are associated with a unique combination of small wavelengths and low frequencies; it is a significant result that a microwave signal can be localized to these phase boundaries by some band topology manipulation.